OpenStudy (anonymous):

if raymond is 12 yrs old and his brother 40. how many years must pass for raymond to be 8 yrs. older than 1/4 as old as his brother?

OpenStudy (ajprincess):

Let x be the number of yrs that shud pass. 12+x=1/4(40+x)+8 12+x=(40+x+(8*4))/4 4(12+x)=40+x+32 4*12+4*x=40+x+32 48+4x=72+x Can u find x @punyeta?

OpenStudy (anonymous):

thank you.